A force 𝐹⃗ = (6πš€Μ‚ - πš₯Μ‚) N acts on a particle that undergoes a displacement 𝑑⃗ = (3πš€Μ‚ + πš₯Μ‚) m. Find (a) the dot product 𝐹⃗ β€’ 𝑑⃗ and (b) the angle between 𝐹⃗ and 𝑑⃗.Β 


1
Expert's answer
2022-01-04T13:54:38-0500

(a) The dot product:


"\vec F·\vec d=(6·3)+((-1)·1)=17\text{ J}."


(b) The angle:


"\theta=\arccos\dfrac{\vec F·\vec d}{|\vec F|·|\vec d|}=28°."